Jennifer has 2 rectangular tablecloths that are similar. The dimensions of the larger tablecloth are 1.5 times as large as the dimensions of the smaller tablecloth. If the perimeter of the larger tablecloth is 24 feet, what is the perimeter of the smaller tablecloth?
Answer by edjones(8007)

Let x & y be 2 adjacent sides of the smaller
p=2x+2y
.
for the larger the sides are 1.5x and 1.5y
P=3x+3y
3x+3y=24
x+y=8
.
3x+3y
2x+2y
------subtract
x+y the difference in perimeters
.
24-8=16 perimeter of the smaller.
.
Ed